OpenStudy (anonymous):

I am valuing a US company with a substantial revenue stream from Korea and Japan. I want to calculate the cost of equity. Some key details for the company - 45% of revenue from Korea and 31% from Japan. According to data on your website country risk premiums for both the countries are 6.28% and 5.75% respectively. Assuming that Koreans companies generate roughly 70% revenue from Korea and Japanese companies have 60% revenue coming from Japan. Assuming 5% implied risk premium for developed markets Can you please tell me the cost of equity equation given this data.

OpenStudy (anonymous):

I think you should use not Total risk premiums for Korea and Japan, but Country risk premiums that are 1.2% and 0.75% respectively. We also need: 1.) Risk free rate. We can assume it is 3.4% 2.) Beta for your company. 3.) Lambda for Korea [% of revenue from Korea] / [1- (Korea’s Export/ Korea’s GDP)] =[45%] / [1-49.9%] = 0.8982 4.) Lambda for Japan [% of revenue from Japan] / [1-(Japan's Export/ Japan's GDP)] =[31%] / [1-12.5%]=0.3543 So, cost of equity is 3.4% + (Beta * 5%) + (Lambda for Korea * 1.2%) + (Lambda for Japan * 0.75%) = 0.034% + (Beta*0.05) + (0.8982*0.012)+(0.3543*0.0075)